OpenStudy (anonymous):

whats the square root of 121

OpenStudy (bahrom7893):

why don't you use a calculator?

OpenStudy (bahrom7893):

Or ask google.. To be honest you should just memorize the squares up to 20.

OpenStudy (anonymous):

11

OpenStudy (bahrom7893):

Seems like too much of an effort to login, and then post a question to just ask something like this.

OpenStudy (anonymous):

\[\sqrt{121} = \sqrt{11 \times 11} = 11\]